OpenStudy (anonymous):

the ratio of the lengths of the corresponding sides of two rectangles is 8:3. The area of the larger rectangle is 320 ft2. What is the area of the smaller rectangle?

Directrix (directrix):

If two polygons are similar, the square of the scale factor of the two polygons is equal to the ratio of any two corresponding area measurements of the polygons.

Directrix (directrix):

(8/3)² = 320/A where A is the area of the smaller polygon 64/9 = 320/A Solve for A.